how to solve (s^5/2)(s^-4/3)(s^-1/6)
Answer by lwsshak3(11628)

how to solve (s^5/2)(s^-4/3)(s^-1/6)
**
(s^5/2)(s^-4/3)(s^-1/6)
change exponents with a common denominator
=(s^15/6)(s^-8/6)(s^-1/6)
add exponents
=s^(15/6-9/6)=s^6/6=s